OpenStudy (anonymous):

Use Function Composition to show the f(x) and g(x) are inverses of each other. y=2x-2 g(x)=2x-2

OpenStudy (anonymous):

they aren't

OpenStudy (anonymous):

the inverse of \[g(x)=2x-2\] is \[g^{-1}(x)=\frac{x+2}{2}\]

OpenStudy (anonymous):

I put the wrong equation. lol. Are \[f(x)=\frac{ 1 }{ 2 }x +1\] and \[g(x)=2x-2\] inveres of each other? @satellite73

OpenStudy (anonymous):

I think they are.